# Quasinormal modes and bound states of massive scalar fields in generalized Bronnikov-Ellis wormholes

### 05 Mar 2024 v2

@@ IV.2 Bound States
In this section we calculate the bound states of massive scalar fields around the generalized Bronnikov-Ellis wormhole using the CFM, which are possible due to the effective potential allows potential wells, see Fig. (1). It is worth to mention that the boundary conditions are different to that of QNMs. In this case we must considering evanescent waves at spatial infinity; therefore, the radial solution that satisfied the boundary condition at infinity is given by
$R\sim e^{-\Omega r}r^{M\omega^{2}/\Omega}\,\,\,\text{as}\,\,\,r\rightarrow\infty\,,$ (29)
where $\Omega=\sqrt{m^{2}-\omega^{2}}$. Generally, in a quantum system with a potential well, there are only a few bound states, the number of which depends on the potential parameters. In several cases, there exists only a single bound state. Thus, we just consider the first bound state which is described by the symmetric solution. Therefore, we consider the following ansatz for the solution to the radial equation (21) which satisfies the desired boundaries conditions
$R(r)=e^{-\Omega r}r^{M\omega^{2}/\Omega}\sum_{n=0}^{\infty}a_{n}\left(\frac{r-b_{0}}{r}\right)^{n}\,.$ (30)
@@
$c_{1}(n)$ $=$ $b_{0}\Bigg{(}\omega^{2}\left(l^{2}+l-4Mn\sqrt{m^{2}-\omega^{2}}-M\sqrt{m^{2}-\omega^{2}}+9n^{2}+1\right)-m^{2}\bigg{(}l^{2}+l-8Mn\sqrt{m^{2}-\omega^{2}}-2M\sqrt{m^{2}-\omega^{2}}$
$+9n^{2}+1\bigg{)}\Bigg{)}-b_{0}^{2}\left(m^{2}-\omega^{2}\right)\left((4n+1)\sqrt{(m-\omega)(m+\omega)}-2m^{2}M\right)+b_{0}^{3}\left(-\left(m^{2}-\omega^{2}\right)^{2}\right)+2M\bigg{(}(l^{2}+l$
@@
$c_{2}(n)$ $=$ $2b_{0}\Bigg{(}m^{2}\left(l^{2}+l+2n\left(-7M\sqrt{m^{2}-\omega^{2}}+4n-6\right)+9M\sqrt{m^{2}-\omega^{2}}+4M^{2}\omega^{2}+6\right)-\omega^{2}\bigg{(}l^{2}+l+$
@@
$\sqrt{(m-\omega)(m+\omega)}-6m^{2}M\bigg{)}+2b_{0}^{3}\left(m^{2}-\omega^{2}\right)^{2}+M\bigg{(}-(6l(l+1)+25n(2n-3)+37)\left(m^{2}-\omega^{2}\right)$
@@
$c_{3}(n)$ $=$ $b_{0}\Bigg{(}\omega^{2}\left(l^{2}+l+2n\left(-9M\sqrt{m^{2}-\omega^{2}}+7n-21\right)+15M\left(2\sqrt{m^{2}-\omega^{2}}+M\omega^{2}\right)+34\right)-m^{2}\bigg{(}l^{2}+l$
@@
$\sqrt{(m-\omega)(m+\omega)}\bigg{)}+b_{0}^{3}\left(-\left(m^{2}-\omega^{2}\right)^{2}\right)+M\bigg{(}3(2l(l+1)+20(n-3)n+49)\left(m^{2}-\omega^{2}\right)-$
@@
$M\left(-(2l(l+1)+20n(2n-9)+209)\left(m^{2}-\omega^{2}\right)+56Mn\omega^{2}\sqrt{m^{2}-\omega^{2}}-121M\omega^{2}\sqrt{m^{2}-\omega^{2}}-18M^{2}\omega^{4}\right)\,,$
@@
and the continued fraction must be truncated at some large index $N$. The bound states frequencies are obtained solving this equation numerically. In Table 1, we show some values of the lowest bound states frequencies $\omega b_{0}$ for scalar perturbations using the CFM. We observe that the oscillation frequency increases when the mass parameter $mb_{0}$ increases and when the angular number $\ell$ increases.
Table 1: The lowest bound states frequencies $\omega b_{0}$ for scalar perturbations for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.25$, $\ell=0,1$, and several values of the mass of the scalar field using the CFM.
@@ V.1 Analytical QNFs of f(R) gravity wormhole
$\Psi(t,r,\theta,\phi)=e^{-i\omega t}\>\frac{R(r)}{r}\>Y_{\ell}(\Omega)\,,$ (34)
@@
$\frac{\mathrm{d}^{2}R}{\mathrm{d}r^{\ast}{}^{2}}+[\omega^{2}-V(r^{\ast})]R=0\,,$ (35)
@@
which we plot in Fig. 6. We observe that the effective potential is symmetric about the throat of the wormhole and takes the form of a barrier for $\ell\neq 0$, with the barrier magnitude increasing with $\ell$. However, it does not allow wells, consequently precluding the support for bound states.
@@
It is worth noticing that when $\ell=0$ the effective potential is constant. Consequently, the solution of the radial function is given by $R=C_{1}e^{i\sqrt{\omega^{2}-\mu^{2}-m^{2}}r^{*}}+C_{2}e^{-i\sqrt{\omega^{2}-\mu^{2}-m^{2}}r^{*}}$. Imposing the condition of purely outgoing waves at both infinities leads to the conclusion that $C_{1}=C_{2}=0$, resulting in $R=0$ for this particular case.
@@
$(1-\tilde{r}^{2})\frac{d^{2}}{d\tilde{r}^{2}}R(\tilde{r})-\tilde{r}\frac{d}{d\tilde{r}}R(\tilde{r})+\left(\frac{\ell(\ell+1)}{\tilde{r}^{2}}-\tilde{\omega}^{2}+1+\tilde{m}^{2}\right)R(\tilde{r})=0\,,$ (38)
@@
$R(\tilde{r})$ $=$ $C_{1}\tilde{r}^{\frac{1}{2}(1-\sqrt{1-4\ell(\ell+1)})}\,_{2}F_{1}\Bigg{(}\frac{1}{4}-\frac{1}{4}\sqrt{1-4\ell(\ell+1)}-\frac{\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}}{2},\frac{1}{4}-\frac{1}{4}\sqrt{1-4\ell(\ell+1)}+\frac{\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}}{2},$ (39)
@@
Since the potential is symmetric about the throat of the wormhole, which is located at $r_{0}^{\ast}=0$, the solutions will be symmetric or anti-symmetric. Therefore, we impose the following boundary conditions at the throat, $\frac{dR}{dr^{\ast}}\left|{}_{r_{0}^{\ast}}=0\right.$ for the symmetric solutions and $R(r_{0}^{\ast})=0$ for the anti-symmetric solutions. These boundary conditions yields the even and odd overtones, respectively. Similar boundary conditions have been applied for other wormhole geometries, where the QNMs were obtained by direct integration of the wave equation Aneesh:2018hlp ; DuttaRoy:2019hij . So, for the symmetric solutions we impose the condition at the throat $\frac{dR}{dr^{\ast}}\left|{}_{r_{0}^{\ast}}\right.=\frac{dR}{d\tilde{r}}\sqrt{1-\frac{b(\tilde{r})}{\tilde{r}}}\left|{}_{\tilde{r}=1}\right.=0$ , which yields
@@
$R(\tilde{r})\sim$ $C_{2}\tilde{r}^{-i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}\Gamma\left[-\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}\right]\Gamma\left[1+\frac{1}{2}\sqrt{1-4\ell(\ell+1)}\right]\Bigg{(}(-1)^{-\alpha_{+}}\rho(\alpha_{-}/\gamma_{-})\frac{\Gamma[1+\alpha_{+}]}{\Gamma[\beta_{-}]\Gamma[1+\gamma_{+}]}$ (40)
@@
Now, in accordance with the boundary condition of purely outgoing waves at spatial infinity, i.e. $R\sim r^{i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}$, we set
@@
On the other hand, we impose the condition at the throat $R(1)=0$ as a boundary condition for the antisymmetric solutions, leading to the expression
@@
$R(\tilde{r})\sim$ $C_{2}\tilde{r}^{-i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}\Gamma\left[-\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}\right]\Gamma\left[1+\frac{1}{2}\sqrt{1-4\ell(\ell+1)}\right]\Bigg{(}-(-1)^{-\alpha_{+}}\frac{\Gamma[\beta_{+}]}{\Gamma[\alpha_{-}]\Gamma[\delta_{+}]}$ (43)
@@
Now, in accordance with the boundary condition of purely outgoing waves at spatial infinity, i.e. $R\sim r^{i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}$, we set
@@
where $\eta^{\prime\prime}\equiv\frac{1}{2}(1+2\ell(\ell+1)-2n(1+n))+\tilde{m}^{2}$ and $n=0,1,\dots$. In Figs. 7 and 8 we plot the behaviour of $Re(\tilde{\omega})$, and $-Im(\tilde{\omega})$ as a function of $\tilde{m}$ by using the analytical expression (46) for $n=0$ and $n=1$, respectively. We can observe that the modes with the lowest angular number $\ell$ exhibit the longest lifetimes, both for the fundamental and the first overtone, and the decay rate decreases with $\tilde{m}$. Additionally, the frequency of oscillation increases with higher values of the parameter $\tilde{m}$ and the angular momentum $\ell$.
@@ V.2 WKB Method
In this section, we employ the Wentzel-Kramers-Brillouin (WKB) approximation method, first introduced by Mashhoon Mashhoon and later developed by Schutz and Iyer Schutz:1985km . This approach allows us to gain insights into the behaviour of QNFs in the eikonal limit as $\ell\rightarrow\infty$. Our aim is to validate and complement the analytical results obtained in the previous section using an alternative method. Iyer and Will computed the third order correction Iyer:1986np , which was subsequently extended to the sixth order Konoplya:2003ii , and recently up to the 13th order Matyjasek:2017psv , see also Konoplya:2019hlu .
@@
This result, as previously mentioned, coincides with the earlier finding, as shown in Eq. (47). Now, we plot the behaviour of $Re(\omega)/\mu$, and $-Im(\omega/\mu)$ as a function of $m/\mu$ by using the 6th order WKB method for $n=0$ (see, Fig. 9), and $n=1$ (see, Fig. 10). We can observe that the modes with the lowest angular number $\ell$ exhibit the longest lifetimes, both for the fundamental and the first overtone. Additionally, the frequency of oscillation increases with higer values of the parameter $m/\mu$ and angular momentum $\ell$. So in this wormhole geometry the anomalous decay of QNFs is avoided.
@@ VI Conclusions
In this work, we explored generalized Bronnikov-Ellis wormholes as background geometries and investigated the propagation of massive scalar fields. The first wormhole spacetime arises as a solution to an action comprising the pure Einstein-Hilbert term and a scalar field with negative kinetic energy. The second wormhole spacetime, on the other hand, results from $f(R)$-gravity.
For the first wormhole spacetime, our analysis focused primarily on low values of the angular number $\ell$, and we determined the QNFs using the CFM. These QNFs exhibit an inverted behaviour of $Im(\omega)b_{0}$: when $mb_{0}>m_{c}b_{0}$, $Im(\omega)b_{0}$ increases with $\ell$, representing an anomalous decay rate of the QNMs. Conversely, for $mb_{0}<m_{c}b_{0}$, $Im(\omega)b_{0}$ decreases with increasing $\ell$ for $n\geq\ell$. The critical value $m_{c}b_{0}$ rises with an increase in the parameter $M/b_{0}$, and the frequency of the oscillation increases with both the angular number and the parameter $mb_{0}$. Additionally, we identified bound states characterized by an oscillation frequency without decay that rises with increasing $mb_{0}$ and $\ell$.
For the second wormhole spacetime, we obtained exact results for the QNFs, which, for large values of $\ell$, coincide with the third-order WKB approximation. Here, the longest-lived modes are those with the lowest angular number $\ell$, both for the fundamental and the first overtone. Furthermore, the oscillation frequency increases with a rise in the parameter $m/\mu$ and angular momentum $\ell$. Thus, in this wormhole background, the anomalous behaviour observed in the first spacetime is avoided. Also, the effective potential is symmetric about the throat of the wormhole and it does not allow wells, consequently bound states are not supported in this wormhole spacetime.

### 10 Apr 2024 v3

@@ IV.2 Bound States
In this section we utilize the CFM to compute the bound states of massive scalar fields around the wormhole geometries defined in (13). These bound states are made feasible by the effective potential, which permits the existence of potential wells for $M\neq 0$, see Fig. (1). It is worth to mention that the boundary conditions are different to that of QNMs. In this case we must considering evanescent waves at spatial infinity; therefore, the radial solution that satisfied the boundary condition at infinity is given by
$u\sim e^{-\Omega r}r^{M\omega^{2}/\Omega}\,\,\,\text{as}\,\,\,r\rightarrow\infty\,,$ (30)
where $\Omega=\sqrt{m^{2}-\omega^{2}}$. Generally, in a quantum system with a potential well, there are only a few bound states, the number of which depends on the potential parameters. In several cases, there exists only a single bound state. Thus, we just consider the first bound state which is described by the symmetric solution. Therefore, we consider the following ansatz for the solution to the radial equation (IV) which satisfies the desired boundaries conditions
$u(r)=e^{-\Omega r}r^{M\omega^{2}/\Omega}\sum_{n=0}^{\infty}a_{n}\left(\frac{r-b_{0}}{r}\right)^{n}\,.$ (31)
@@
$c_{1}(n)$ $=$ $b_{0}\Bigg{(}\omega^{2}\left(\ell^{2}+\ell-4Mn\sqrt{m^{2}-\omega^{2}}-M\sqrt{m^{2}-\omega^{2}}+9n^{2}+1\right)-m^{2}\bigg{(}\ell^{2}+\ell-8Mn\sqrt{m^{2}-\omega^{2}}-2M\sqrt{m^{2}-\omega^{2}}$
$+9n^{2}+1\bigg{)}\Bigg{)}-b_{0}^{2}\left(m^{2}-\omega^{2}\right)\left((4n+1)\sqrt{(m-\omega)(m+\omega)}-2m^{2}M\right)+b_{0}^{3}\left(-\left(m^{2}-\omega^{2}\right)^{2}\right)+2M\bigg{(}(\ell^{2}+\ell$
@@
$c_{2}(n)$ $=$ $2b_{0}\Bigg{(}m^{2}\left(\ell^{2}+\ell+2n\left(-7M\sqrt{m^{2}-\omega^{2}}+4n-6\right)+9M\sqrt{m^{2}-\omega^{2}}+4M^{2}\omega^{2}+6\right)-\omega^{2}\bigg{(}\ell^{2}+\ell+$
@@
$\sqrt{(m-\omega)(m+\omega)}-6m^{2}M\bigg{)}+2b_{0}^{3}\left(m^{2}-\omega^{2}\right)^{2}+M\bigg{(}-(6\ell(\ell+1)+25n(2n-3)+37)\left(m^{2}-\omega^{2}\right)$
@@
$c_{3}(n)$ $=$ $b_{0}\Bigg{(}\omega^{2}\left(\ell^{2}+\ell+2n\left(-9M\sqrt{m^{2}-\omega^{2}}+7n-21\right)+15M\left(2\sqrt{m^{2}-\omega^{2}}+M\omega^{2}\right)+34\right)-m^{2}\bigg{(}\ell^{2}+\ell$
@@
$\sqrt{(m-\omega)(m+\omega)}\bigg{)}+b_{0}^{3}\left(-\left(m^{2}-\omega^{2}\right)^{2}\right)+M\bigg{(}3(2\ell(\ell+1)+20(n-3)n+49)\left(m^{2}-\omega^{2}\right)-$
@@
$M\left(-(2\ell(\ell+1)+20n(2n-9)+209)\left(m^{2}-\omega^{2}\right)+56Mn\omega^{2}\sqrt{m^{2}-\omega^{2}}-121M\omega^{2}\sqrt{m^{2}-\omega^{2}}-18M^{2}\omega^{4}\right)\,,$
@@
and the continued fraction must be truncated at some large index $N$. The bound states frequencies are obtained solving this equation numerically. In Tables 1 and 2, we show some values of the lowest bound states frequencies $\omega b_{0}$ for scalar perturbations using the CFM. We observe that the oscillation frequency increases when the mass parameter $mb_{0}$ increases and when the angular number $\ell$ increases.
Table 1: Frequencies ($\omega b_{0}$) of the lowest bound states for scalar perturbations in the wormhole spacetime described by Eq. (13), with $M/b_{0}=0.1$, $\ell=0,1$, and several values of the scalar field mass using the CFM.
$\ell$ $mb_{0}=0.4$ $mb_{0}=0.8$ $mb_{0}=1.2$ $mb_{0}=1.6$
0 0.39967670 0.79727718 1.18979435 1.57334204
1 0.39991988 0.79935626 1.19781061 1.59474718
$\ell$ $mb_{0}=2$ $mb_{0}=2.4$ $mb_{0}=2.8$ $mb_{0}=3.2$
0 1.94709369 2.31390013 2.67665885 3.03714485
1 1.98954025 2.38132146 2.76858541 3.14873064
Table 2: Frequencies ($\omega b_{0}$) of the lowest bound states for scalar perturbations in the wormhole spacetime described by Eq. (13), with $M/b_{0}=0.25$, $\ell=0,1$, and several values of the scalar field mass using the CFM.
@@ V.1 Analytical QNFs of f(R) gravity wormhole
$\Psi(t,r,\theta,\varphi)=e^{-i\omega t}\>\frac{u(r)}{r}\>Y_{\ell}(\Omega)\,,$ (35)
@@
$\frac{\mathrm{d}^{2}u}{\mathrm{d}r^{\ast}{}^{2}}+[\omega^{2}-V(r^{\ast})]u=0\,,$ (36)
@@
which is plotted in Fig. 8. We observe that the effective potential is symmetric about the throat of the wormhole and takes the form of a barrier for $\ell\neq 0$, with the barrier magnitude increasing with $\ell$. However, it does not allow wells, consequently precluding the support for bound states.
@@
It is worth noticing that when $\ell=0$ the effective potential is constant. Consequently, the solution of the radial function is given by $u=C_{1}e^{i\sqrt{\omega^{2}-\mu^{2}-m^{2}}r^{*}}+C_{2}e^{-i\sqrt{\omega^{2}-\mu^{2}-m^{2}}r^{*}}$. Imposing the condition of purely outgoing waves at both infinities leads to the conclusion that $C_{1}=C_{2}=0$, resulting in $u=0$ for this particular case.
@@
$(1-\tilde{r}^{2})\frac{d^{2}}{d\tilde{r}^{2}}u(\tilde{r})-\tilde{r}\frac{d}{d\tilde{r}}u(\tilde{r})+\left(\frac{\ell(\ell+1)}{\tilde{r}^{2}}-\tilde{\omega}^{2}+1+\tilde{m}^{2}\right)u(\tilde{r})=0\,,$ (39)
@@
$u(\tilde{r})$ $=$ $C_{1}\tilde{r}^{\frac{1}{2}(1-\sqrt{1-4\ell(\ell+1)})}\,_{2}F_{1}\Bigg{(}\frac{1}{4}-\frac{1}{4}\sqrt{1-4\ell(\ell+1)}-\frac{\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}}{2},\frac{1}{4}-\frac{1}{4}\sqrt{1-4\ell(\ell+1)}+\frac{\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}}{2},$ (40)
@@
Since the potential is symmetric about the throat of the wormhole, which is located at $r_{0}^{\ast}=0$, the solutions will be symmetric or anti-symmetric. Therefore, we impose the following boundary conditions at the throat, $\frac{du}{dr^{\ast}}\left|{}_{r_{0}^{\ast}}=0\right.$ for the symmetric solutions and $u(r_{0}^{\ast})=0$ for the anti-symmetric solutions. These boundary conditions yields the even and odd overtones, respectively. For the symmetric solutions the condition at the throat $\frac{du}{dr^{\ast}}\left|{}_{r_{0}^{\ast}}\right.=\frac{du}{d\tilde{r}}\sqrt{1-\frac{b(\tilde{r})}{\tilde{r}}}\left|{}_{\tilde{r}=1}\right.=0$ yields
@@
$u(\tilde{r})\sim$ $C_{2}\tilde{r}^{-i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}\Gamma\left[-\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}\right]\Gamma\left[1+\frac{1}{2}\sqrt{1-4\ell(\ell+1)}\right]\Bigg{(}(-1)^{-\alpha_{+}}\rho(\alpha_{-}/\gamma_{-})\frac{\Gamma[1+\alpha_{+}]}{\Gamma[\beta_{-}]\Gamma[1+\gamma_{+}]}$ (41)
@@
Now, in accordance with the boundary condition of purely outgoing waves at spatial infinity, i.e. $u\sim r^{i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}$, we set
@@
On the other hand, the boundary condition at the throat $u(1)=0$ for the antisymmetric solutions leads to the expression
@@
$u(\tilde{r})\sim$ $C_{2}\tilde{r}^{-i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}\Gamma\left[-\sqrt{1+\tilde{m}^{2}-\tilde{\omega}^{2}}\right]\Gamma\left[1+\frac{1}{2}\sqrt{1-4\ell(\ell+1)}\right]\Bigg{(}-(-1)^{-\alpha_{+}}\frac{\Gamma[\beta_{+}]}{\Gamma[\alpha_{-}]\Gamma[\delta_{+}]}$ (44)
@@
Now, in accordance with the boundary condition of purely outgoing waves at spatial infinity, i.e. $u\sim r^{i\sqrt{\tilde{\omega}^{2}-1-\tilde{m}^{2}}}$, we set
@@
where $\eta^{\prime\prime}\equiv\frac{1}{2}(1+2\ell(\ell+1)-2n(1+n))+\tilde{m}^{2}$ and $n=0,1,\dots$. In Figs. 9 and 10 we plot the behaviour of $Re(\tilde{\omega})$, and $-Im(\tilde{\omega})$ as a function of $\tilde{m}$ by using the analytical expression (47) for $n=0$ and $n=1$, respectively. We can observe that the modes with the lowest angular number $\ell$ exhibit the longest lifetimes, both for the fundamental and the first overtone, and the decay rate decreases with $\tilde{m}$. Additionally, the frequency of oscillation increases with higher values of the parameter $\tilde{m}$ and the angular momentum $\ell$.
@@ V.2 WKB Method
In this section, we employ the Wentzel-Kramers-Brillouin (WKB) approximation method, first introduced by Mashhoon Mashhoon and later developed by Schutz and Iyer Schutz:1985km , to obtain the QNFs. This approach allows us to gain insights into the behaviour of QNFs in the eikonal limit as $\ell\rightarrow\infty$. Our aim is to validate and complement the analytical results obtained in the previous section using an alternative method. Iyer and Will computed the third order correction Iyer:1986np , which was subsequently extended to the sixth order Konoplya:2003ii , and recently up to the 13th order Matyjasek:2017psv , see also Konoplya:2019hlu .
@@
This result, as previously mentioned, coincides with the earlier finding, as shown in Eq. (48). Now, we plot the behaviour of $Re(\omega)/\mu$, and $-Im(\omega/\mu)$ as a function of $m/\mu$ by using the 6th order WKB method for $n=0$ (see, Fig. 11), and $n=1$ (see, Fig. 12). We can observe that the modes with the lowest angular number $\ell$ exhibit the longest lifetimes, both for the fundamental and the first overtone. Additionally, the frequency of oscillation increases with higer values of the parameter $m/\mu$ and angular momentum $\ell$. Therefore, in this wormhole geometry the anomalous decay of QNFs is avoided.
@@ VI Conclusions
In this work, we explored wormhole geometries with a non-constant redshift function by introducing a gravitational mass $M$, which goes over into the Bronnikov-Ellis wormhole when the gravitational mass parameter vanishes and an $f(R)$ gravity wormhole spacetime as background geometries and investigated the propagation of massive scalar fields.
For the wormholes with a non-constant reshift function, our analysis focused primarily on low values of the angular number $\ell$, and we determined the QNFs using the CFM. These QNFs exhibit an inverted behaviour of $Im(\omega)b_{0}$: when $mb_{0}>m_{c}b_{0}$, $Im(\omega)b_{0}$ increases with $\ell$, representing an anomalous decay rate of the QNMs. Conversely, for $mb_{0}<m_{c}b_{0}$, $Im(\omega)b_{0}$ decreases with increasing $\ell$ for $n\geq\ell$. The critical value $m_{c}b_{0}$ rises with an increase in the parameter $M/b_{0}$, and the frequency of the oscillation increases with both the angular number and the parameter $mb_{0}$. Additionally, for the geometries with $M\neq 0$ we identified bound states characterized by an oscillation frequency without decay that rises with increasing $mb_{0}$ and $\ell$. On the other hand, for $M=0$ the effective potential does not forms wells, and bound states are prohibited in this case.
Then, we considered a wormhole spacetime that is an exact solution of $f(R)$ modified gravity with a phantom scalar field with a self-interacting potential. We obtained exact results for the QNFs, which, for large values of $\ell$, coincide with the third-order WKB approximation. Here, the longest-lived modes are those with the lowest angular number $\ell$, both for the fundamental and the first overtone. Furthermore, the oscillation frequency increases with a rise in the parameter $m/\mu$ and angular momentum $\ell$. Thus, in this wormhole background, the anomalous behaviour observed in the first spacetimes is avoided. Also, the effective potential is symmetric about the throat of the wormhole and it does not allow wells, consequently bound states are not supported in this wormhole spacetime.
The analysis of the wormholes considered reveals that those with a constant redshift function do not support bound states, in contrast to those with a non-constant redshift function (resulting in a non-zero tidal radial force), which do permit bound states.